OpenStudy (anonymous):

Please help I have no ide what this is -14=-4(9x-1)

OpenStudy (anonymous):

x=1/2

OpenStudy (anonymous):

distribute

OpenStudy (anonymous):

-36x + 4 = -14 -36x = -18 x= 1/2